Nina read 2/16 of her book Wuthering Heights. Her brother Juan had read 4/48 of his book, a tale of two cities. How much more had Nina read than Juan?
1
Expert's answer
2020-02-24T11:13:31-0500

Solution. Nina read 2/16=1/8 of her book.  Her brother Juan had read 4/48=1/12 of his book.

Find how much more had Nina read than Juan.


"\frac {1}{8}- \frac{1}{12}=\frac{3-2}{24}=\frac{1}{24}"

Answer.

"\frac{1}{24}"